Shawn and Dorian rented bikes from two different rental shops. The prices in dollars, y, of renting bikes from the two different shops for x hours is shown.
Shop Shawn used: y = 10 + 3.5x
Shop Dorian used: y = 6x
If Shawn and Dorian each rented bikes for the same number of hours and each paid the same price, how much did each pay for the rental?

Answer by jorel555(1290)

10+3.5x=6x
2.5x=10
x=10/2.5=4 hours
They each paid 24!!!!!!!!